OpenStudy (anonymous):

identify the equation as a conditional equation, a contradiction , or an identity . then describe the solution can anyone explain these to me? 7x+3=6(x-2)

OpenStudy (anonymous):

Solve the equation for x first. 7x+3 = 6x -12 -6x x + 3 = -12 -3 x = -15 A contradiction is something that is never true. Something like -1=1. An identity is something that is always true. Something like 1=1. Is x = -15 always true? Never true? Or is it conditional on x?

OpenStudy (anonymous):

What does it mean by describing the solution?

OpenStudy (anonymous):

so it is conditional?

OpenStudy (anonymous):

yes - its conditional because it is only true for some or one value of x

OpenStudy (anonymous):

jimmy caan u help me with a couple more?

OpenStudy (anonymous):

I can.